# VizGroup: An AI-Assisted Event-Driven System for Real-Time Collaborative Programming Learning Analytics

### v1

@@ Abstract.
Programming instructors often conduct collaborative learning activities, like Peer Instruction, to foster a deeper understanding in students and enhance their engagement with learning. These activities, however, may not always yield productive outcomes due to the diversity of student mental models and their ineffective collaboration. In this work, we introduce VizGroup, an AI-assisted system that enables programming instructors to easily oversee students’ real-time collaborative learning behaviors during large programming courses. VizGroup leverages Large Language Models (LLMs) to recommend event specifications for instructors so that they can simultaneously track and receive alerts about key correlation patterns between various collaboration metrics and ongoing coding tasks. We evaluated VizGroup with 12 instructors using a dataset collected from a Peer Instruction activity that was conducted in a large programming lecture. The results showed that compared to a version of VizGroup without the suggested units, VizGroup with suggested units helped instructors create additional monitoring units on previously undetected patterns on their own, covered a more diverse range of metrics, and influenced the participants’ following notification creation strategies.
@@ 1. Introduction
To assess VizGroup’s usability and effectiveness, we conducted a between-subject study with 12 participants with teaching experience. Participants used a basic visual analytics tool without a notification system and then used VizGroup with or without our LLM recommendation notifications. The results showed that compared to a version of VizGroup without the notification recommendation, VizGroup with suggested units helped instructors create additional monitoring units that were previously undiscoverable on their own. These recommendation covered a more diverse range of metrics, providing a more comprehensive understanding of the learning process. Furthermore, we found evidence that the suggested notifications influenced the participants’ decision-making when selecting the following monitoring unit criteria.
@@ 2.3. Collaboration Analytics
Prior work, such as Pair-Up (Xu et al., 2023), explored collaboration analytics by examining the transition behaviors of K-12 students from individual to group learning. Groupdynamics (Sato et al., 2023) provided a summary of the vocal activities and statuses of up to 10 small discussion groups. While these systems provided valuable insights into collaboration dynamics, they focused on prescribed metrics and lacked support for discovering the emerging correlation patterns between collaboration and the tasks that people are working on.
@@ 2.4. Notification Systems
Notification is an important feature in today’s software systems. It allows users to receive relevant updates more timely. Common examples include monitors and alert functions in medical systems that are designed to support doctors in receiving time-critical decision support (Mastrianni et al., 2022a, b), and email and message notifications in mobile applications that are designed to help users stay aware of personal matters (Li et al., 2023). However, these works often focus on the impact that notifications have on users’ work and well-being, such as work disruption, and the appropriate time to receive notifications (Auda et al., 2018; Ho et al., 2018). In contrast, we look at how to help users choose and create notifications to optimize their resources or effort in real-time.
@@ 3. Formative Study
Due to the lack of real-time collaborative learning VA systems, we deployed a technology probe (Hutchinson et al., 2003) to explore the potential informational needs of instructors and the corresponding design challenges. To discover design considerations and derive interface designs for our probe, we first conducted semi-structured interviews with four experienced instructors of large programming courses at our institution. These need-finding interviews focused on the types of information instructions desired in real-time during the group discussion activity and the decisions they aimed to make based on this information. We found that instructors desired to discern patterns in code and group discussions and their interplay in how discussions aided students in resolving their issues.
@@ 3.2. Methodology
We recruited 8 participants with experience teaching programming courses to participate in our within-subject study where they evaluated the probe and a baseline system. Participants then completed a survey and interview. We captured an in-class peer instruction session from a large-scale introductory university programming course (i.e., 111 students, 37 groups) and used its live playback during the study. The baseline condition was an ablated version of the probe without the group discussion visualization and intelligent features like discussion topic summary, team activity level, and message tags, similar to existing educational VA systems.
@@
Figure 2. VizGroup’s User Interface. (A) Scatterplot displays student’s progress via group view, group structure view and individual view. (B) Suggested notification panel that displays context-aware notification suggestions and user-defined notifications. (C) Active notification panel that displays active trackers and alerts. (D) Group information panel that shows group and individual activity. (E) LLM summary of student conversation and conversation history. (F) Class performance that shows the number of students that passes each unit tests. (G) List of Details on group conversation topic and student error messages. (H) Each session can be controlled with playback.
@@ 3.3. Results and Design Considerations
To analyze the results, we calculated the correctness of the identified trend (i.e., a particular student’s or group’s task pass rate and discussion increased or decreased over time), and the precision and recall of the classification tasks (i.e., identify all the students that have engaged in group discussions over a period of time). For Task 1, a T-test found no significant difference in the precision ($p=0.11$) and recall ($p=0.44$) between the two conditions. For Task 2, participants had a higher recall using VizGroup (VizGroup: $Mean=0.990,\sigma=0.029$; Baseline: $Mean=0.742,\sigma=0.200,p<0.01$), while the precision was on par for both conditions ($p=0.96$). We also found that participants had better accuracy while tracking and understanding group dynamics trends during Task 3 while using the probe ($Mean=0.875,\sigma=0.173$) than the baseline system ($Mean=0.667,\sigma=0.178,p<0.05$). We did not find a significant difference in the time spent completing the tasks (VizGroup: $Mean=1277.29s$, Baseline: $Mean=1244.86s$, $p>0.05$).
@@
From the post-study survey and interviews, we found that the probe’s ability to present discussion topics and activity levels prompted participants to investigate patterns starting from group engagement down to student-level coding challenges. Moreover, the probe’s high-level information display and visualization features streamlined the process of monitoring and analyzing collaborative learning dynamics, thereby facilitating a more nuanced understanding of student and group performance over time. However, the probe demanded higher cognitive loads as participants needed to manually keep track of an identified pattern over time, making the observation unscalable to monitor multiple patterns (e.g., students who were not engaging, students who lacked support, etc.). In addition, different patterns were relevant in different contexts. For example, initially, students may not be engaged in a discussion as they work on a coding task, but this idleness could become an issue once everyone else has submitted their code. The context dependency prompts a design that would enable contextually-aware pattern discovery and monitoring in the next iteration of our system design.
@@ 4. Design Goals
Based on these findings and the key challenges identified in prior work, three design goals guided the iterative development of VizGroup to support instructors in being able to easily monitor collaborative learning in large programming classes.
@@
Conversation Topic Each group’s topic summary based on their existing chat messages. Individual conversation topics is tagged by LLM (Figure  1.4b) but is not used in visualization and notifications. Varies, example topics include Correcting Function Logic and Syntax Errors, Troubleshooting Code and EOF Errors
@@
Figure 3. Overview of collaborative learning visualization with Group, Structure, and Individual Views, switchable by zooming. Features: (1) History trace of data points upon selection, (2) Area highlight for inspecting multiple points, (3) Structure View plots participants against conversation topics, (4) Group data point location changes with topic or participation shifts, (5) Arrows indicate participation levels, (6) Flashing effect for changes in pass rates or interactions.
  • DG1: Enable instructors to efficiently navigate and analyze collaborative learning dynamics at different levels of granularity, from class-wide overviews to detailed views of individual student performance.
  • DG2: Enable instructors to easily create customizable trackers and alerts to monitor desired patterns and temporal changes in real-time, while minimizing interference with their workflow.
  • DG3: Provide context-sensitive notifications that use a pattern’s urgency and relevance to an instructor’s current context as indicators of when an alert should be provided.
5. VizGroup
As VizGroup expanded on the probe VA system, it shared similar features including (1) a chat panel that utilized the LLM to analyze group and individual activity (Figures  2C and  1.3), an LLM summary of group discussion topics (Figures  2D and  1.4a,b) and a class performance panel (Figures 2E and  1.2). We also modified the detailed information table to display an aggregated list of conversation topics when the Group view was selected on the scatterplot or an aggregated list of student code errors when the Individual view was selected on the scatterplot.
During a class session, VizGroup recorded data on an individual student and on a group level (Table  1). On an individual level, VizGroup tracked each student’s pass rate, activity level, and type of code errors. On the group activity level, it tracked the average pass rate of the group, the activity level of the group, the group’s conversation topic, and the team structure of the group.
5.1. Three Level View of Collaborative Learning
Figure 4. Clicking on a group avatar displays its information summary panel, and activity and performance table
5.1.1. Group View:
This view presents an aggregated perspective of the group’s collaborative interactions, enabling instructors to gauge the average group activity level(y-axis) and average group pass rate (x-axis) over time (Figure  3). By aggregating data over collective activities, this view facilitates an understanding of group dynamics and collaborative patterns at a glance.
@@ 5.1.2. Structure View:
This view maps the current topics of discussion (x-axis) against the number of active members in a group (y-axis), offering a detailed snapshot of group dynamics (Figure  3.3). We implemented a topic modeling approach to assess the relevance of each chat message and to distill the overarching subjects of discussion. Initially, we utilized a LLM to generate text embeddings for each conversation, which serve as a nuanced representation of the discussions’ content. These embeddings were then clustered using the K-means++ algorithm, a method chosen for its efficiency in grouping data. For each identified cluster, we produced a concise summary that captures the essence of the conversations within. To ensure a streamlined and coherent compilation of conversation topics, we compared new summaries with existing ones. If a newly generated summary closely matched any of the previously established summaries, we classified it under that pre-existing category.
We integrated graph representations and incorporated indicators of activity levels directly within the view to simplify the analysis process for instructors. As the activity levels or the discussion topics change, each group cluster data point will transition to a different position of the visualization (Figure  3.4). Each group’s communication pattern is also depicted using arrows to indicate the sender and the recipient of chat messages, with the arrow’s thickness representing the level of activity amongst group members (e.g. no messages [Uncaptioned image], few messages [Uncaptioned image],many messages [Uncaptioned image]; Figure  3.5). The thickness of the arrow is calculated by:
@@
where A is the activity level of the individual. The green dots indicating that students passed all unit tests, whereas the pink indicated they did not fully pass all unit tests. When there were updates to the group conversation topics or the number of active members, a flashing animation was applied to the changed data point to catch the instructor’s attention (Figure  3.6). To prevent information overload, VizGroup displayed up to eight group structures with the lowest pass rate simultaneously. Expanding the lowest performing groups would enable instructors to identify if group activity contributed to the low pass rates.
@@ 5.1.3. Individual View:
This view enables instructors to assess each student’s engagement (y-axis) and performance (x-axis) closely (Figure  3). It was designed to highlight individual behaviors, such as students who might be struggling or feeling excluded, in relation to their peers’ progress. This view aids instructors in quickly identifying students in need of additional support or those who are excelling and can potentially serve as resources for their peers.
Figure 6. Alerts notify instructors if students/groups meets the defined criteria. (1) Count of data points/groups that matches specified criteria, (2) Alert type (group or individual), (3) Notification creation reason, (4) Trigger criteria (modifiable via highlighted area click), (5) Criteria selection drop-down, (6) Clickable current data points meeting criteria, displayed on scatter plot and detail panels.(7) Instructors can change between using a spatial threshold or a temporal threshold. (8) Instructors can preview data points on the scatter plot. (9) An alert activation confirmation is sent to the list of Active Notifications. (10) Instructors can select the attributes to be included in the threshold. (11) Flashing Animation when alert is triggered.
@@ 5.1.4. Interaction
To support seamless navigation across these views (DG1), VizGroup enables instructors to zoom in and out. Selecting a region (Figure  3.1) or clicking on a data point provides further insights in the detailed information list and the group details panel (Figure  4). Clicking on a data point will also show the trace history of the selected data point, which enables instructors to view changes in performance over time (Figure  3.2).
@@ 5.2. Notifications for Observing Student Activity
Notifications are displayed as two columns: Suggested Notifications (Figure  2B) and Active Notifications (Figure  2C). Instructors can add and edit existing trackers or alerts in the Suggested Notification panel before activating it. After activation, the notification will appear in the Active Notifications panel, where it will start tracking real time data and push alerts based on instructor-specified criteria.
5.2.1. Trackers (Figure  5)
Trackers are visualizations that display the current count of student activity attributes. Instructors can select a variable to be tracked by clicking on the highlighted attribute, which will reveal three options: Code Issues, Conversation Topics, and Members Participated (Figure  5.1). They can display the visualization as a bar chart that shows the count of each values within a specific attribute (e.g., Code Issues are grouped by the count of different errors, Conversation Topics are grouped by the count of different groups of conversation summaries whereas Members Participated is grouped the the total number of active participants in the group). They can also view a time-series line plot that displays the changes in the different groups of the selected attribute from the start of the session until the current time (Figure  5). Instructors can toggle between the two visualization easily (Figure  5.2).

### v2

@@ Abstract.
Programming instructors often conduct collaborative learning activities, like Peer Instruction, to foster a deeper understanding in students and enhance their engagement with learning. These activities, however, may not always yield productive outcomes due to the diversity of student mental models and their ineffective collaboration. In this work, we introduce VizGroup, an AI-assisted system that enables programming instructors to easily oversee students’ real-time collaborative learning behaviors during large programming courses. VizGroup leverages Large Language Models (LLMs) to recommend event specifications for instructors so that they can simultaneously track and receive alerts about key correlation patterns between various collaboration metrics and ongoing coding tasks. We evaluated VizGroup with 12 instructors in a comparison study using a dataset collected from a Peer Instruction activity that was conducted in a large programming lecture. The results showed that VizGroup helped instructors effectively overview, narrow down, and track nuances throughout students’ behaviors.
Programming Education, Collaborative Learning
journalyear: 2024copyright: acmlicensedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676347isbn: 979-8-4007-0628-8/24/10
@@ 1. Introduction
To assess VizGroup’s usability and effectiveness, we conducted a between-subject study with 12 participants with teaching experience. Participants used a basic visual analytics tool without a notification system and then used VizGroup with or without our LLM recommendation notifications. The results showed that compared to a version of VizGroup without the notification recommendation, VizGroup with suggested units helped instructors create additional monitoring units that were previously undiscoverable on their own. These recommendations covered a more diverse range of metrics, providing a more comprehensive understanding of the learning process. Furthermore, we found evidence that the suggested notifications influenced the participants’ decision-making when selecting the following monitoring unit criteria.
@@ 2.3. Collaboration Analytics
Prior work, such as Pair-Up (Yang et al., 2023), explored collaboration analytics by examining the transition behaviors of K-12 students from individual to group learning. Groupdynamics (Sato et al., 2023) provided a summary of the vocal activities and statuses of up to 10 small discussion groups. While these systems provided valuable insights into collaboration dynamics, they focused on prescribed metrics and lacked support for discovering the emerging correlation patterns between collaboration and the tasks that people are working on.
@@ 2.4. Notification Systems
Notification is an important feature in today’s software systems. It allows users to receive relevant updates more timely. Common examples include monitors and alert functions in medical systems that are designed to support doctors in receiving time-critical decision support (Mastrianni et al., 2022a, b), and email and message notifications in mobile applications that are designed to help users stay aware of personal matters (Li et al., 2023). There were also works on providing real-time notifications and feedback in classrooms to enhance teachers’ awareness (Martinez-Maldonado et al., 2014) and support teachers’ orchestration (Tissenbaum and Slotta, 2019). However, these works often focus on the impact that notifications have on users’ work and well-being, such as work disruption, and the appropriate time to receive notifications (Auda et al., 2018; Ho et al., 2018). In contrast, we look at how to help users choose and create notifications to optimize their resources or effort in real-time.
@@ 3. Formative Study
Due to the lack of real-time large-scale collaborative learning VA systems, we deployed a technology probe (Hutchinson et al., 2003) to explore the potential informational needs of instructors and the corresponding design challenges. To discover design considerations and derive interface designs for our probe, we first conducted semi-structured interviews with four experienced instructors of large programming courses at our institution. These need-finding interviews focused on the types of information instructions desired in real-time during the group discussion activity and the decisions they aimed to make based on this information. We found that instructors desired to discern patterns in code and group discussions and their interplay in how discussions aided students in resolving their issues.
@@ 3.2. Methodology
We recruited 8 participants with experience teaching programming courses to participate in our within-subject study where they evaluated the probe and a baseline system. Participants then completed a survey and interview. We captured an in-class peer instruction session from a large-scale introductory university programming course (i.e., 111 students, 37 groups) and used its live playback during the study. The baseline condition was an ablated version of the probe without the group discussion visualization and intelligent features like discussion topic summary, team activity level, and message tags (i.e., without 4a, 4b, 5a, and 5b in Figure 2), similar to existing educational VA systems.
@@
Figure 3. VizGroup’s User Interface. (1) Scatterplot displays student’s progress via group view, group structure view and individual view. (2) Suggested notification panel that displays context-aware notification suggestions and user-defined notifications. (3) Active notification panel that displays active trackers and alerts. (4) Class performance that shows the number of students that passes each unit tests. (5) List of Details on group conversation topic and student error messages. (6) Group information panel that shows group and individual activity. (7) LLM summary of student conversation and conversation history. (8) Each session can be controlled with playback.
@@ 3.3. Results and Design Considerations
To analyze the results, we calculated the correctness of the identified trend (i.e., a particular student’s or group’s task pass rate and discussion increased or decreased over time), and the precision and recall of the classification tasks (i.e., identify all the students that have engaged in group discussions over a period of time). For Task 1, a T-test found no significant difference in the precision ($p=0.11$) and recall ($p=0.44$) between the two conditions. For Task 2, participants had a higher recall using the probe (Probe: $Mean=0.990,\sigma=0.029$; Baseline: $Mean=0.742,\sigma=0.200,p<0.01$), while the precision was on par for both conditions ($p=0.96$). We also found that participants had better accuracy while tracking and understanding group dynamics trends during Task 3 while using the probe ($Mean=0.875,\sigma=0.173$) than the baseline system ($Mean=0.667,\sigma=0.178,p<0.05$). We did not find a significant difference in the time spent completing the tasks (Probe: $Mean=1277.29s$, Baseline: $Mean=1244.86s$, $p>0.05$).
@@
From the post-study survey and interviews, we found that the probe’s ability to present discussion topics and activity levels prompted participants to investigate patterns starting from group engagement down to student-level coding challenges. 6 participants (P1, P3, P5, P6, P7, P8) reported discussion topic as one of the ”most satisfying” features because it is ”very useful for instructor to find the common issues” without “checking it one by one” (P6). P4 also suggested that activity levels addressed the issue of “a little bit nervous (seeing) automatically generated topics”. Moreover, participants reported that the probe’s high-level information display and visualization features “provided about what was going on in the class.” (P5) and “allow me to quickly overview the status of the students, and help me easily identify the student who is struggling” (P2). These features streamlined the process of monitoring and analyzing collaborative learning dynamics, thereby facilitating a more nuanced understanding of student and group performance over time.
However, the probe demanded higher cognitive loads as participants needed to manually keep track of an identified pattern over time, making the observation unscalable to monitor multiple patterns (e.g., students who were not engaging, students who lacked support, etc.). P6 stated that “sometimes I forget some of the students’ performance before that I can’t tell if the pass rate was increased.” Participants also mentioned the importance of focusing on a “level or window” when inspecting students and groups “that can make it more efficient as we are not distracted by other information” (P5). In addition, different patterns were relevant in different contexts. For example, initially, students may not be engaged in a discussion as they work on a coding task, but this idleness could become an issue once everyone else has submitted their code. The context dependency prompts a design that would enable contextually-aware pattern discovery and monitoring in the next iteration of our system design.
@@ 4. Design Goals
Our formative study revealed two primary effects driving our DGs: (1) instructors used different levels of information scope to understand student progress, and (2) instructors were overloaded when tracking progress and providing support. Based on these findings and the key challenges identified in prior work, three design goals guided the iterative development of VizGroup to support instructors in being able to easily monitor collaborative learning in large programming classes.
  • DG1. Efficient and contextualized navigation across different levels of granularity. To enable instructors to identify patterns of concern in real-time using interactive learning analytics, the navigation process must be both efficient and adaptable to different levels of granularity, allowing seamless transitions between individual, group, and class-wide perspectives.
  • DG2. Real-time and attention-free monitoring. To address the overload instructors experienced when trying to follow the progress of parallel groups, there is a need for a system that offers real-time, attention-free monitoring, along with mechanisms to selectively guide their attention and minimize interference with their workflow.
  • DG3. Guiding instructors’ attention while considering both instructors’ and students’ needs in context. To avoid compounding instructors’ overload with unnecessary interruptions, mechanisms for guiding instructors’ attention (e.g., via notifications) must be context-aware, taking into account the needs of both instructors and students.
Figure 4. Overview of collaborative learning visualization with Group, Structure, and Individual Views, switchable by zooming. Features: (1) Area highlight for inspecting multiple points, (2) History trace of data points upon selection, (3) Structure View plots participants against conversation topics, (4) Group data point location changes with topic or participation shifts, (5) Arrows indicate participation levels, (6) Flashing effect for changes in pass rates or interactions.
5. VizGroup
As VizGroup expanded on the probe VA system, it shared similar features including (1) a chat panel that utilized the LLM to analyze group and individual activity (Figure  3.7 and  2.3), an LLM summary of group discussion topics (Figure  3.6 and  2.4a,b) and a class performance panel (Figure  3.4 and  2.2). We also modified the detailed information table to display an aggregated list of conversation topics when the Group view was selected on the scatterplot or an aggregated list of student code errors when the Individual view was selected on the scatterplot.
During a class session, VizGroup recorded data on an individual student and on a group level (Table  1). On an individual level, VizGroup tracked each student’s pass rate, activity level, and type of code errors. On the group activity level, it tracked the average pass rate of the group, the activity level of the group, the group’s conversation topic, and the team structure of the group. We chose these five dimensions because of their critical importance, as highlighted by previous research (Yang et al., 2023; Sato et al., 2023; Zhang et al., 2024), and based on the preferences and experience of instructors from the formative study and the data we have. Pass Rate serves as a proxy for students’ progress in understanding and utilizing the learning objectives, while Activity Level indicates students’ engagement in collaboration. Likewise, Team Structure reflects students’ participation at a group level. Code Issue highlights errors as identified by the compiler in their submissions, and Conversation Topic reveals what students encounter in their collaborations.
5.1. Three Level View of Collaborative Learning
Figure 5. Clicking on a group data point displays its information summary panel, and activity and performance table.
5.1.1. Group View:
This view presents an aggregated perspective of the group’s collaborative interactions, enabling instructors to gauge the average group activity level (y-axis) and average group pass rate (x-axis) over time (Figure  4). By aggregating data over collective activities, this view facilitates an understanding of group dynamics and collaborative patterns at a glance.
@@ 5.1.1. Group View:
Conversation Topic Each group’s topic summary based on their existing chat messages. Individual conversation topics is tagged by LLM (Figure  2.4b) but is not used in visualization and notifications. Varies, example topics include Correcting Function Logic and Syntax Errors, Troubleshooting Code and EOF Errors
@@ 5.1.2. Structure View:
This view maps the current topics of discussion (x-axis) against the number of active members in a group (y-axis), offering a detailed snapshot of group dynamics (Figure  4.3). We implemented a topic modeling approach to assess the relevance of each chat message and to distill the overarching subjects of discussion. Initially, we utilized an LLM to generate text embeddings for each conversation, which serve as a nuanced representation of the discussions’ content. These embeddings were then clustered using the K-means++ (Bachem et al., 2016) algorithm, a method chosen for its efficiency and reliability in grouping semantic word representations (Ding et al., 2024). For each identified cluster, we produced a concise summary that captures the essence of the conversations within. To ensure a streamlined and coherent compilation of conversation topics, we compared new summaries with existing ones. If a newly generated summary closely matched any of the previously established summaries, we classified it under that pre-existing category.
We integrated graph representations and incorporated indicators of activity levels directly within the view to simplify the analysis process for instructors. As the activity levels or the discussion topics change, each group cluster data point will transition to a different position of the visualization (Figure  4.4). Each group’s communication pattern is also depicted using arrows to indicate the sender and the recipient of chat messages, with the arrow’s thickness representing the level of activity amongst group members (e.g. no messages [Uncaptioned image], few messages [Uncaptioned image],many messages [Uncaptioned image]; Figure  4.5). The thickness of the arrow is calculated by:
@@
where A is the activity level of the individual. The green dots indicating that students passed all unit tests, whereas the pink indicated they did not fully pass all unit tests. When there were updates to the group conversation topics or the number of active members, a flashing animation was applied to the changed data point to catch the instructor’s attention (Figure  4.6). To prevent information overload, VizGroup displayed up to eight group structures with the lowest pass rate simultaneously. Expanding the lowest performing groups would enable instructors to identify if group activity contributed to the low pass rates.
@@ 5.1.3. Individual View:
This view enables instructors to assess each student’s engagement (y-axis) and performance (x-axis) closely (Figure  4). It was designed to highlight individual behaviors, such as students who might be struggling or feeling excluded, in relation to their peers’ progress. This view aids instructors in quickly identifying students in need of additional support or those who are excelling and can potentially serve as resources for their peers.
Figure 7. Alerts notify instructors if students/groups meets the defined criteria. (1) Count of data points/groups that matches specified criteria, (2) Alert type (group or individual), (3) Notification creation reason, (4) Trigger criteria (modifiable via highlighted area click), (5) Criteria selection drop-down menu, (6) Clickable current data points meeting criteria, displayed on scatter plot and detail panels.(7) Instructors can change between using a spatial threshold or a temporal threshold. (8) Instructors can preview data points on the scatter plot. (9) An alert activation confirmation is sent to the list of Active Notifications. (10) Instructors can select the attributes to be included in the threshold. (11) Flashing Animation when alert is triggered.
@@ 5.1.4. Interaction
To support seamless navigation across these views (DG1), VizGroup enables instructors to zoom in and out. Selecting a region (Figure  4.1) or clicking on a data point provides further insights in the detailed information list and the group details panel (Figure  5). Clicking on a data point will also show the trace history of the selected data point, which enables instructors to view changes in performance over time (Figure  4.2).
@@ 5.2. Notifications for Observing Student Activity
Notifications are displayed as two columns: Suggested Notifications (Figure  3.2) and Active Notifications (Figure  3.3).Instructors can add and edit existing trackers or alerts in the Suggested Notification panel before activating it. After activation, the notification will appear in the Active Notifications panel, where it will start tracking real time data and push alerts based on instructor-specified criteria.
@@
5.2.1. Trackers (Figure  6)
Trackers are visualizations that display the current count of student activity attributes. Instructors can select a variable to be tracked by clicking on the highlighted attribute, which will reveal three options: Code Issues, Conversation Topics, and Members Participated (Figure  6.1). They can display the visualization as a bar chart that shows the count of each values within a specific attribute (e.g., Code Issues are grouped by the count of different errors, Conversation Topics are grouped by the count of different groups of conversation summaries whereas Members Participated is grouped the the total number of active participants in the group). They can also view a time-series line plot that displays the changes in the different groups of the selected attribute from the start of the session until the current time (Figure  6). Instructors can toggle between the two visualization easily (Figure  6.2).